OpenStudy (anonymous):

Three number cubes are tossed, what is the probability of two number cubes showing sis, if the first cube showed 6

OpenStudy (anonymous):

so you have two more throws left, and you must get one mor ehead.

OpenStudy (anonymous):

So you want to get one more six, and then one more of anything but a six.

OpenStudy (anonymous):

= P( six, not six) + P(not six, six) = (1/6)(5/6) x 2

OpenStudy (anonymous):

= (1/3)